In how many ways can we select a chairperson, vice- chairperson, and recorder from a group of 11 persons?


1
Expert's answer
2021-11-05T06:03:54-0400

Let the first selected person would be a chairperson, second - a vice-chairperson, and third - a recorder. Then we have to select 3 people fom 11, and the order of the selection is important(for example, A, B, C woulb be different from B, A, C). So, the point is to calculate the number of permutations when repetition is not allowed

"P(11, 3) = {\frac {11!} {(11-3)!}}=11*10*9=990" ways